Zuo Software categorizes its accounts receivable into four age groups for purposes of estimating its allowance for uncollectible accounts.

  1. Accounts not yet due = $385,000; estimated uncollectible = 15%.

  2. Accounts 1–30 days past due = $48,200; estimated uncollectible = 18%.

  3. Accounts 31–90 days past due = $38,500; estimated uncollectible = 22%

  4. Accounts more than 90 days past due = $28,800; estimated uncollectible = 25%.

At 12/31/2021, before recording any adjustments, Zuo has a credit balance of $25,300 in its allowance for uncollectible accounts.

Required:
1. Estimate the appropriate 12/31/2021 balance for Zuo’s allowance for uncollectible accounts.
2. What journal entry should Zuo record to adjust its allowance for uncollectible accounts?
3. Calculate Zuo’s 12/31/2021 net accounts receivable balance.
Answer #1

Req 1)

385,000*15%=57,750
48,200*18%=8,676
38,500*22%=8,470
28,800*25%=7,200

balance in allowance for 

uncollectible accounts(answer)

82,096


Req 2)

82,096-25,300= 56,796

dr. bad debt expense    56,769

 cr. allowance for uncollectible accounts  56,769


Req 3)

385,000+48,200+38,500+28,800= 500,500

500,500- 82,096= 418,404 net accounts receivable balance (answer)

answered by: anonymous